A four-year project promises to pay 0.56% interest per month. An investor is trying to decide between investing in this project or investing in a semi-annually compounded four-year GIC. Above what interest rate would the GIC need to offer to make it a more appealing investment?
Answer #1

Suppose Investor ready to invest  $10,000.

Assume A is Value of $10,000 invested for 04 years in project equals to,

P = Present Value of Investment = 10,000

i = Monthly interest rate 0.56% = 0.56/100 = 0.0056

n = Period = 4 *12 = 48 Months (As interest paid monthly)

So,

= 10,000 * 1.3074 = 13074.13

Now to make GIC Project a more appealing Investor must have return higher than 13074.13.

So,

A = 13074.13

P =10,000

i = semi annual compounded interest rate

n = Period = 4*2 = 8 As interest rate paid Semi annualy

So,

%

So to make GIC more appealing to the investor it should offer more than 3.40% interest rate (Semi Annual)